HENDI SPECTROSCOPY OF C-H STRETCHING MODES OF METHANE AND SUBSTITUTED METHANE MOLECULES

\underline{ROBERT R. FEHNEL} and KEVIN K. LEHMANN, Department of Chemistry, University of Virginia, Charlottesville VA, 22904-4319.

We report the first results with a new Helium Nanodroplet Isolation Spectroscopy machine at UVa. The excitation source is an ARGOS cw PPLN OPO (Aculight). which produces ∼2 W in the 3.2-3.9 μm spectral window. We are extending the spectroscopy of methane in helium, having observed weak, Fermi Resonance induced ro-vibrational transitions of CH_{4} and the C-H fundamentals of CH_{3}D, CH_{2}D_{2}, and CHD_{3}. We have also observed the spectrum of CH_{3}Cl and plan to report on other substituted methane molecules at the conference.